OpenStudy (anonymous):

The sun always illuminates half of the moon’s surface, except during a lunar eclipse. The illuminated portion of the moon visible from Earth varies as it revolves around Earth resulting in the phases of the moon. The period from a full moon to a new moon and back to a full moon is called a synodic month and is 29 days, 12 hours, and 44.05 minutes long. Write a sine function that models the fraction of the moon’s surface which is seen to be illuminated during a synodic month as a function of the number of days, d, after a full moon. [Note: full moon equals 1/2 illuminated.] I'm lost.

OpenStudy (whpalmer4):

Okay, you need to make a sine function with period 29 days, 12 hours, and 44.05 minutes (you'll need to express that as a decimal number of days, which will be slightly greater than 29.5). Then you'll want to make sure the phase of your sine function is correct so that at t = 0 and at t = 29 days 12 hours 44.05 minutes the value of your sine function reaches its maximum, which should be 1/2 (at full moon, half the surface is illuminated). So, you'll have something like \[F = \frac{1}{2}\sin (kt + p)\] where k is chosen to give you kt = the period of an ordinary sine function when t = 29 days 12 hours 44.05 minutes, and p is whatever fudge factor you need to add to make the sin function = 1 at t = 0. I fear I didn't help with that explanation :-)

OpenStudy (anonymous):

Okai I got the first part, but the second and third part you just completely lost me

OpenStudy (whpalmer4):

Well, let's break it down. What is the period of the sine function \(y=\sin t\)?

OpenStudy (anonymous):

29=sin1/2?

OpenStudy (whpalmer4):

no, ignore the moon part of the problem, we'll just do straight trig functions. what is the period of sine function \(y = \sin \theta\)?

OpenStudy (whpalmer4):

If we start at \(\theta = 0\), what is the value of \(\sin\theta\) there?

OpenStudy (whpalmer4):

What is \(\sin(0)\)?

OpenStudy (anonymous):

sin(29) = .4848096202?

OpenStudy (anonymous):

and sin 0 is 0

OpenStudy (whpalmer4):

Please, forget the moon problem, just answer the questions I ask, thanks. Okay, \(sin(0) = 0\). What is \(\sin(\pi/2)\)?

OpenStudy (anonymous):

0.274121336

OpenStudy (anonymous):

And sorry (._. )

OpenStudy (whpalmer4):

mmm...no. Are you computing in degrees or radians? Radians is what we want here. \(\sin(\pi/2) = 1\) — that's pointing straight up to the top of the unit circle.

OpenStudy (whpalmer4):

right? I drew in the radius at \(\theta = \pi/2\)

OpenStudy (anonymous):

oh my bad I have it set on degrees

OpenStudy (anonymous):

okai I got 1 now

OpenStudy (anonymous):

Oh I'm suppose to be using the unit circle in this one?

OpenStudy (whpalmer4):

We're just exploring the properties of the sine function, so we can figure out how to bend it to our will :-)

OpenStudy (anonymous):

So how would I set this up?

OpenStudy (whpalmer4):

So, the period of the sine function (just like the cosine function) is \(2\pi\) radians. Every \(2\pi\) radians the radius makes a complete trip around the unit circle: